Given that x^n is inversely proportional to y and y is a quarter of its original value when x is doubled, calculate the value of n.
Answer #1

We have xn = k/y where k is the constant of proportionality. Hence y = k/xn.

Aso y/4 = k/(2x)n so that y = 4k/(2n *xn).

Therefore, k/xn = 4k/(2n *xn) or, (2n *xn)/ xn = = 4k/k or, 2n = 4 = 22.

Hence n = 2.